The sound from a trumpet radiates uniformly in all directions in air with a temperature of 20?C. At a distance of 4.86m from the trumpet the sound intensity level is 53.0dB . The frequency is 585Hz

A)What is the pressure amplitude at this distance?
Use 344m/s for the speed of sound in air, 1.20kg/m3 for the density of air, and 1.00
Answer #1

c)

sound intensity level, beta = 10*log(I1/IO)

53 = 10*log(I1/10^-12)

5.3 = log(I/10^-12)

10^5.3 = I/10^-12

I1 = 1.99*10^-7 W/m^2

sound intensity level, beta = 10*log(I2/IO)

26 = 10*log(I2/10^-12)

2.6 = log(I2/10^-12)

10^2.6 = I2/10^-12

I2 = 3.98*10^-10 W/m^2

we know, I = P/(4*pi*r^2


so, I1/I2 = (r2/r1)^2

r2 = r1*sqrt(I1/I2)

= 4.86*sqrt(1.99*10^-7/3.98*10^-10)

= 108.7 m